OpenStudy (anonymous):

@satellite73 Solve the system of equations by addition 3x + 6y = -15 4x - 6y = -6

OpenStudy (anonymous):

could have sworn i just answered this one add in a column, get \[7x=-21\] solve for x in one step, what do you get?

OpenStudy (anonymous):

oops! I meant to put in x + 2y = 1 5x + 4y = -7

OpenStudy (anonymous):

multiply the first one by -2 all the way across, get \[-2x-4y=1\\ 5x+4y=-7\] then add in a column and the y terms drop out

OpenStudy (anonymous):

let me know when you get \[3x=-6\]

OpenStudy (anonymous):

(-3,2)